OpenStudy (anonymous):

Find a formula for the derivatives of g(x) = 2x^2-3 using difference quotients?

zepdrix (zepdrix):

So they want you to use the definition of the derivative :D which is kind of a pain in the butt... XD remember that one or no? :D

OpenStudy (anonymous):

yea f(x+h) -f(x) / h

zepdrix (zepdrix):

yah, so you take the "limit" of the difference quotient, letting h approach 0. and that will give you an equation of the line tangent to g(x) at any point. (in other words, it will give you the derivative).. setting up the problem is the tricky part :) need help with that?

OpenStudy (anonymous):

yea thats where I am screwing up, with all the algebra and putting the equ together

zepdrix (zepdrix):

\[g(x)=2x^2 -3 \rightarrow g(x+h)=2(x+h)^2 -3\]

zepdrix (zepdrix):

\[\frac{ g(x+h)-g(x) }{ h } = \frac{ 2(x+h)^2-3 -[2x^2-3] }{ h }\]

zepdrix (zepdrix):

setup look ok? c:

OpenStudy (anonymous):

yep I got this far

zepdrix (zepdrix):

expanding out the binomial will give us..\[\frac{ 2x^2+4xh+2h^2-3-2x^2+3 }{ h }\]

OpenStudy (anonymous):

ok I can't expand the binomial correctly

zepdrix (zepdrix):

ok ignore the 2 for a moment :) so its the square of the first term + 2 times the product of the terms in the middle + the square of the 2nd term \[(x+h)^2=(x+h)(x+h)=x^2+xh+xh+h^2=x^2+2xh+h^2\]

OpenStudy (anonymous):

I wasn't ignoring the 2. :)

OpenStudy (anonymous):

wait ok I need to see how to put the 2 back into the equ

zepdrix (zepdrix):

\[2(x+h)^2=2(x^2+2xh+h^2)\]

OpenStudy (anonymous):

what is this called so I can get specific practice on this stuff, just factoring?

zepdrix (zepdrix):

i think it's called "expanding binomials". its a binomial since it has two terms. and you multiply them out, expanding or something XD

OpenStudy (anonymous):

ok cool

OpenStudy (anonymous):

ok then can I factor a 2 from the numerator?

zepdrix (zepdrix):

if you get time, you should check out pascal's triangle. there is a really important connection between binomials and his silly triangle :D it makes it much much easier to expand 3rd degree binomials and higher

zepdrix (zepdrix):

i dont htink you want to factor out a 2 D: since you have some nice cancellations

OpenStudy (anonymous):

oops I forgot about the rest of the equ ok Pascals triangle , got it thanks!

OpenStudy (anonymous):

ok I get 4xh+2h^2 /h

zepdrix (zepdrix):

ah yes sounds good so far! :)

OpenStudy (anonymous):

got disconnected from the web

OpenStudy (anonymous):

ok so factor 2h from top and get 4x+h? the book answer is 4x

zepdrix (zepdrix):

ok good, you got to answer of 4x+h, ONLY NOW should you take the limit, allow h to approach 0. we don't run into problems anymore since we're no longer dividing by h.